### v1

# Positivity bounds from thermal field theory entropy

Abstract
We present a novel approach to deriving positivity bounds on effective field theories by analyzing the thermodynamic behavior of thermal quantum field systems. Focusing on scalar theories with higher-dimensional operators, we compute the finite-temperature entropy using thermal field theory techniques. We argue that consistency with fundamental thermodynamic principles—specifically, the expectation that entropy increases with the introduction of new degrees of freedom—imposes nontrivial constraints on Wilson coefficients. In particular, we show that the coefficient of the leading dimension-8 operator must be strictly positive. This thermodynamic perspective offers an alternative to traditional S-matrix-based derivations of positivity bounds and provides fresh insight into the interplay between entropy, unitarity, and causality in quantum field theory.
I Introduction
Effective Field Theory (EFT) is a powerful framework for describing low-energy phenomena in systems characterized by a hierarchy of energy scales. In particular, EFTs are valuable when the ultraviolet (UV) completion of a theory is either unknown or strongly coupled, such as in gravitational or hadronic systems. Two primary approaches exist for constructing EFTs: the top-down method, which integrates out heavy degrees of freedom from a known UV theory, and the bottom-up method, which enumerates all operators consistent with low-energy symmetries, treating their coefficients as free parameters.
However, if we assume the existence of a consistent UV completion that respects causality, locality, and unitarity, then the low-energy Wilson coefficients are not entirely arbitrary [1]. Constraints arising from these fundamental principles—commonly referred to as positivity bounds—have traditionally been derived using analyticity properties of scattering amplitudes and dispersion relations [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
In this work, we explore an alternative approach. Motivated by previous studies connecting the thermodynamic properties of black hole entropy to quantum gravity conjectures [57, 58, 59, 60, 61, 62, 63, 64, 65, 66], we investigate whether similar principles can be applied within the framework of finite-temperature quantum field theory. Our central proposal is that enforcing thermodynamic consistency—specifically, requiring that the thermal entropy does not decrease when additional degrees of freedom are introduced—naturally leads to positivity constraints on the Wilson coefficients in the EFT expansion.
To make this idea concrete, we analyze a scalar field theory with a shift symmetry, compute its entropy at finite temperature, and show that ensuring a physically meaningful entropy implies a strictly positive coefficient for the leading higher-dimensional operator. This suggests a deep connection between statistical mechanics and the consistency structure of effective field theories, offering a complementary perspective to traditional amplitude-based arguments.
II Basic idea
Consider a system in thermal equilibrium characterized by two widely separated mass scales: a heavy mass $M$ and a light mass $m$. We examine the system at an intermediate temperature T satisfying $m\ll T\ll M$, where the heavy modes are Boltzmann suppressed and can be integrated out, while the light modes remain thermally active. In this regime, the full thermal entropy of the system can be approximated using an effective field theory that includes only the light degrees of freedom:
$S_{m+M}\simeq S_{m\text{-EFT}}.$ (1)
Naively, one expects that introducing new heavy degrees of freedom should only increase the thermal entropy, as more microscopic configurations become accessible. This intuition naturally leads to the inequality:
$S_{m+M}>S_{m-\text{only}}.$ (2)
At the classical level, it is natural to expect that the entropy of the full system $m+M$ should exceed that of the light sector $m$ alone. Motivated by this inequality, one might be tempted to compute the thermal entropy within the effective field theory and compare it with that of the free field theory, anticipating that such a comparison would yield a corresponding positivity condition.
However, this classical intuition does not always hold in quantum theory, as the situation becomes more subtle when quantum entanglement is involved. For example, consider a combined system $A+B$ prepared in a pure quantum state, so that the total entropy vanishes, $S_{A+B}=0$. If we only observe subsystem $A$ by tracing out $B$, the reduced state of $A$ generally becomes mixed.
In this case, although the total entropy is zero, the entropy of the observed subsystem can still be nonzero due to quantum entanglement:
$S_{A+B}=0\quad\text{but}\quad S_{A}>0.$ (3)
Therefore, in addition to the thermal entropy of $m$ and $M$, we must try to include the entanglement entropy between them. That is,
$S_{m}=S_{m,\text{thermal}}+S_{m\rightarrow M,\text{entangled}},$ (4)
and
$S_{M}=S_{M,\text{thermal}}+S_{M\rightarrow m,\text{entangled}}.$ (5)
This entanglement entropy is generally ultraviolet divergent and depends on the choice of regulator, so it does not capture the thermal dynamical information we aim to extract. We will encounter this divergence again when performing thermal calculations. For our purposes, the only property we need to emphasize is that the entanglement entropy between two subsystems is symmetric:
$S_{m\rightarrow M,\text{entangled}}=S_{M\rightarrow m,\text{entangled}}.$ (6)
We now invoke the Araki-Lieb inequality [67, 68] to derive a bound on the total entropy:
$S_{M+m}\,\geq\,\big{|}S_{m}-S_{M}\big{|}.$ (7)
Equality is achieved only when the global state is pure or when one of the subsystems is trivial (i.e., has zero entropy).
Combining Eqs. (1, 4-7), at temperature $T$ satisfying $m\ll T\ll M$, we finally obtain:
$S_{m,\text{EFT}}>|S_{m,\text{thermal}}-S_{M,\text{thermal}}|.$ (8)
At temperatures much lower than the mass of a heavy particle $(T\ll M)$, the high-energy degrees of freedom are difficult to excite. As a result, the thermal entropy is exponentially suppressed: $S_{M,thermal}\propto T^{\frac{3}{2}}e^{-\frac{M}{T}}$, This contribution can be approximated as zero. Therefore, the inequality becomes:
$S_{m,\text{EFT}}>S_{m,\text{thermal}}.$ (9)
This inequality fits our physical intuition (2), but is now rigorously justified after accounting for the contribution from entanglement entropy.
III Calculation of Thermal entropy
We begin by considering a simple scalar field theory that possesses a shift symmetry, which physically corresponds to taking the massless limit $(m\to 0)$. Such a field often emerges as a Goldstone boson in effective field theories, representing the low-energy fluctuations around spontaneously broken continuous symmetries. Throughout this section, we adopt natural units with $k_{B}=\hbar=1$.
The low-energy effective Lagrangian for such a scalar field $\phi$ up to dimension 8 operator is generally given by:
$\mathcal{L}_{M}=\frac{1}{2}\left(\partial\phi\right)^{2}+\frac{c}{M^{4}}\left(\partial\phi\right)^{4}+\cdots.$ (10)
where $c$ is a dimensionless coupling constant, and $M$ denotes a high-energy cutoff scale in the effective theory. The first term corresponds to the standard kinetic energy of a free scalar field, while the second term represents the leading-order self-interactions of the field.
To incorporate finite-temperature effects, we consider the system in thermal equilibrium with a heat bath at temperature $T$, which allows us to compute the statistical properties of the scalar field within a consistent quantum field theoretical framework.
From a theoretical perspective, to prepare a thermal equilibrium state at temperature $T=\tfrac{1}{\beta}$, we can employ the Euclidean path integral formalism on a compact manifold with temporal circumference $\beta$. Concretely, this approach involves the following steps [69]:
  • Imaginary Time (Wick Rotation): We replace real time $t$ by imaginary (Euclidean) time $\tau$ through the transformation:
    $\tau\equiv it.$ (11)
    This step transforms the Minkowski spacetime into a Euclidean spacetime, making the field theory resemble a classical statistical system.
  • Euclidean Lagrangian: Under Wick rotation, the Lagrangian transforms into its Euclidean counterpart. The Euclidean Lagrangian $L_{E}$ is related to the original Minkowski Lagrangian by:
    $L_{E}(\tau,x)\equiv-\mathcal{L}_{M}(t,x),$ (12)
  • Periodic Boundary Conditions and Temperature: At finite temperature, the Euclidean time direction becomes compactified with period $\beta$, where
    $\tau\sim\tau+\beta,\text{ where }\beta\equiv\frac{1}{T}.$ (13)
    This periodicity reflects the thermal equilibrium condition. And this emerges naturally from the trace over states in the definition of the partition function:
    $\mathcal{Z}(T)=\text{Tr}(e^{-\beta H}).$ (14)
Following these steps, the Euclidean action for the scalar theory becomes:
$S_{E}$ $=S_{0}+S_{I}$
$=\int^{\beta}_{0}d\tau\int_{\textbf{x}}\left[\frac{1}{2}\left(\partial\phi\right)^{2}-\frac{c}{M^{4}}\left(\partial\phi\right)^{4}+\cdots.\right]$ (15)
where we clearly separate the free action $S_{0}$ (quadratic in $\phi$) from the interaction terms $S_{I}$ (higher-order terms in $\phi$). Note that in the zero-temperature limit $(\beta\to\infty)$, the periodicity condition disappears, and we naturally recover the standard Euclidean quantum field theory defined on infinite spacetime.
A crucial motivation for studying this theory at finite temperature is understanding how thermal entropy, a fundamental quantity describing the disorder and microscopic states of the system, constrains the properties of the theory. In particular, as we shall see from detailed calculations, consistency conditions related to the entropy demand the positivity of certain coupling constants. In this model, we will see that this implies:
$c>0.$ (16)
Thus, we can find that thermal field theory not only provides a rich physical framework to study quantum systems at finite temperature but also imposes powerful theoretical constraints on allowed interactions.
To quantify these ideas precisely, we construct the thermal partition function:
$\mathcal{Z}(T)=C\int\mathcal{D}\phi e^{-S_{E}}=C\int\mathcal{D}\phi e^{\left(-S_{0}-S_{I}\right)}.$ (17)
Given that the coupling $\frac{c}{M^{4}}$ is small (since $M$ is large compared to other physical scales in the theory), we can expand the partition function perturbatively around the free theory. Explicitly, we obtain:
$\mathcal{Z}(T)$ $=C\int\mathcal{D}\phi e^{-S_{0}}\left[1-S_{\mathrm{I}}+\frac{1}{2}S_{\mathrm{I}}^{2}-\frac{1}{6}S_{\mathrm{I}}^{3}+\cdots\right]$
$=\mathcal{Z}_{(0)}\left[1-\left\langle S_{\mathrm{I}}\right\rangle_{0}+\frac{1}{2}\left\langle S_{\mathrm{I}}^{2}\right\rangle_{0}-\frac{1}{6}\left\langle S_{\mathrm{I}}^{3}\right\rangle_{0}+\cdots\right],$ (18)
where
$\mathcal{Z}_{(0)}\equiv C\int\mathcal{D}\phi e^{-S_{0}},$ (19)
$\langle\cdots\rangle_{0}\equiv\frac{\int\mathcal{D}\phi[\cdots]\exp\left(-S_{0}\right)}{\int\mathcal{D}\phi\exp\left(-S_{0}\right)}.$ (20)
The free energy density $f(T)$ can then be computed from the partition function as follows:
$f(T)\equiv\lim_{V\to\infty}\frac{F(T,V)}{V},\text{ with }F(T,V)=-T\ln\mathcal{Z}.$ (21)
It is standard to separate the free energy into contributions from the free theory and corrections due to interactions
$f=f_{0}+f_{\left(\geq 1\right)}.$ (22)
For leading-order and first-order, we have:
$f_{(0)}$ $=\lim_{V\to\infty}\frac{T}{V}S_{0}=\lim_{V\to\infty}\frac{-T\ln\mathcal{Z}_{0}}{V},$ (23)
$f_{(1)}$ $=\lim_{V\to\infty}\frac{T}{V}\langle S_{I}\rangle_{0}$
$=\lim_{V\to\infty}\!\!-\frac{T}{V}\!\int^{\beta}_{0}\!\!\!d\tau\!\!\int_{\textbf{x}}\!\frac{c}{M^{4}}\!\langle\partial_{\mu}\phi(X)\partial^{\mu}\phi(X)\partial_{\nu}\phi(X)\partial^{\nu}\phi(X)\rangle_{0}.$ (24)
(a) $f_{(0)}$
(b) $f_{(1)}$
Figure 1: Diagrams contributing to our condition.
These two quantities correspond to the FIG. (1). Before carrying out the actual calculations, we can make rough estimates. For $f_{(0)}$ the result involves a single momentum integral and scales as $\sim T^{4}$. For $f_{(1)}$, based on the diagram in FIG. (1(b)), there are two independent momenta and the calculation will be two separated integrations of these two momenta, and dimensional analysis suggests a scaling of the form $\frac{c}{M^{4}}T^{8}$.
We now proceed to compute $\mathcal{Z}_{0}$ from which $f_{(0)}$ can be extracted. The partition function $\mathcal{Z}_{0}$ takes the form:
$\mathcal{Z}_{0}$ $=C\int\mathcal{D}\phi\exp\left[-\int_{0}^{\beta}d\tau\int d^{3}\mathbf{x}\,\frac{1}{2}(\partial\phi)^{2}\right].$ (25)
We adopt the following Fourier expansion for the scalar field:
$\phi\left(X\right)=\sqrt{\beta V}\sum_{\omega_{n}}\int\frac{d^{d}\textbf{k}}{\left(2\pi\right)^{d}}\tilde{\phi}(K)e^{iK\cdot X}\equiv\mathop{\mathrlap{\sum}\int}_{K}\tilde{\phi}(K)e^{iK\cdot X},$ (26)
where $K\equiv(\omega_{n},\textbf{k})$, with bosonic Matsubara frequencies [70] $\omega_{n}=2\pi Tn,\ n\in\mathbb{Z}$, imposed by the periodicity $e^{i\omega_{n}\beta}=1$. Furthermore, for the real scalar field, we have the condition: $\phi^{\ast}(X)=\phi(X)\Rightarrow\tilde{\phi}^{\ast}(K)=\tilde{\phi}(-K)$. Substituting this into the action gives:
$\frac{\mathcal{Z}_{0}}{C}$ $=\int\mathcal{D}\phi\exp\left[\int_{0}^{\beta}d\tau\int d^{3}\mathbf{x}\frac{1}{2}\mathop{\mathrlap{\sum}\int}_{K,K^{\prime}}(K_{\mu}K^{\prime\mu})\right$.
$\left.\tilde{\phi}(K)\tilde{\phi}(K^{\prime})\times e^{iK\cdot X+iK^{\prime}\cdot X}\right]$
$=\int\mathcal{D}\phi\exp\left[-\frac{1}{2}\beta V\sum_{\omega_{n}}\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}K^{2}\tilde{\phi}(-K)\tilde{\phi}(K)\beta\right]$
$=\int\mathcal{D}\phi\exp\left[-\frac{1}{2}\beta^{2}V\sum_{\omega_{n}}\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}K^{2}\tilde{\phi}^{\ast}(K)\tilde{\phi}(K)\right]$
$=\int\mathcal{D}\phi\left[\sum_{K}-\frac{1}{2}\tilde{\phi}^{\ast}(K)K^{2}\beta^{2}\tilde{\phi}(K)\right].$ (27)
[…]
$\mathcal{Z}_{0}=C\cdot\sqrt{\frac{1}{\det(K^{2}\beta^{2})}},$ (28)
where the constant $C$ contributes only to the vacuum energy and can be dropped since it is independent of temperature. Therefore:
$\ln\mathcal{Z}_{0}$ $=-\frac{1}{2}\ln\det(K^{2}\beta^{2}),$
$=-\frac{1}{2}\sum_{K}\ln\left(K^{2}\beta^{2}\right)=-\frac{1}{2}V\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\sum_{\omega_{n}}\ln(K^{2}\beta^{2}),$
$=V\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\left[-\frac{1}{2}|\mathbf{k}|-T\cdot\ln\left(1-e^{-|\mathbf{k}|/T}\right)\right].$ (29)
where the last step uses the standard Matsubara sum formula. Thus, the free energy density becomes:
$\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\left[\frac{1}{2}|\mathbf{k}|+T\cdot\ln\left(1-e^{-|\mathbf{k}|/T}\right)\right].$ (30)
The leading temperature-independent term reflects the zero-temperature vacuum energy, which lacks thermal significance. As such, it can be consistently subtracted when evaluating thermodynamic quantities. This yields the renormalized result:
$f_{(0)}$ $=\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}T\cdot\ln\left(1-e^{-|\mathbf{k}|/T}\right),$
$=\frac{T}{2\pi^{2}}\int dk\cdot k^{2}\ln\left(1-e^{-k/T}\right)$
$=\frac{T^{4}}{2\pi^{2}}\int dx\cdot x^{2}\ln\left(1-e^{-x}\right),$
$=-\frac{\pi^{2}T^{4}}{90}.$ (31)
In the third line, we performed a change of variable $k/T\to x$. Then we turn to the evaluation of $f_{(1)}$, from the last line of Eq. (24), we note that $\langle\partial_{\mu}\phi(X)\partial^{\mu}\phi(X)\partial_{\nu}\phi(X)\partial^{\nu}\phi(X)\rangle_{0}$ is independent of spacetime coordinates due to translational invariance, then the spacetime integral becomes trivial. So the $f_{(1)}$ can be written as
$f_{(1)}=-\frac{c}{M^{4}}\langle\partial_{\mu}\phi(0)\partial^{\mu}\phi(0)\partial_{\nu}\phi(0)\partial^{\nu}\phi(0)\rangle_{0}.$ (32)
To handle the derivatives, we perform a Fourier transform of the fields:
$f_{(1)}$ $=-\frac{c}{M^{4}}\langle\mathop{\mathrlap{\sum}\int}_{K_{1}}\ iK_{1,\mu}\tilde{\phi}\left(K_{1}\right)\mathop{\mathrlap{\sum}\int}_{K_{2}}\ iK_{2}^{\mu}\tilde{\phi}\left(K_{2}\right)$
$\times\mathop{\mathrlap{\sum}\int}_{K_{3}}\ iK_{3,\nu}\tilde{\phi}\left(K_{3}\right)\mathop{\mathrlap{\sum}\int}_{K_{4}}\ iK_{4}^{\nu}\tilde{\phi}\left(K_{4}\right)\rangle_{0}$
$=-\frac{c}{M^{4}}\mathop{\mathrlap{\sum}\int}_{K_{1},K_{2},K_{3},K_{4}}K_{1,\mu}K_{2}^{\mu}K_{3,\nu}K_{4}^{\nu}$
$\times\langle\tilde{\phi}(K_{1})\tilde{\phi}(K_{2})\tilde{\phi}(K_{3})\tilde{\phi}(K_{4})\rangle_{0}.$ (33)
Using Wick’s theorem, we have:
$f_{(1)}=$ $-\frac{c}{M^{4}}\mathop{\mathrlap{\sum}\int}_{K_{1},K_{2},K_{3},K_{4}}K_{1,\mu}K_{2}^{\mu}K_{3,\nu}K_{4}^{\nu}$
$\times\left[\langle\tilde{\phi}(K_{1})\tilde{\phi}(K_{2})\rangle_{0}\langle\tilde{\phi}(K_{3})\tilde{\phi}(K_{4})\rangle_{0}\right$.
$\left.+\langle\tilde{\phi}(K_{1})\tilde{\phi}(K_{3})\rangle_{0}\langle\tilde{\phi}(K_{2})\tilde{\phi}(K_{4})\rangle_{0}\right$.
$\left.+\langle\tilde{\phi}(K_{1})\tilde{\phi}(K_{4})\rangle_{0}\langle\tilde{\phi}(K_{2})\tilde{\phi}(K_{3})\rangle_{0}\right].$ (34)
With $\langle\tilde{\phi}(K_{1})\tilde{\phi}(K_{2})\rangle_{0}=\frac{T^{2}}{V}\delta_{\omega_{1_{n}}+\omega_{2_{n}},0}(2\pi)^{d}\delta^{(d)}(\mathbf{k_{1}}+\mathbf{k_{2}})\frac{1}{K_{1}^{2}}$, we obtain:
$f_{(1)}=$ $-\frac{c}{M^{4}}T^{2}\sum_{\omega_{1_{n}}\omega_{3_{n}}}\int\frac{d^{3}\mathbf{k_{1}}}{(2\pi)^{3}}\frac{d^{3}\mathbf{k_{3}}}{(2\pi)^{3}}K_{1}^{2}K_{3}^{2}\frac{1}{K_{1}^{2}}\frac{1}{K_{3}^{2}}$
$-2\frac{c}{M^{4}}T^{2}\sum_{\omega_{1_{n}}\omega_{2_{n}}}\int\frac{d^{3}\mathbf{k_{1}}}{(2\pi)^{3}}\frac{d^{3}\mathbf{k_{2}}}{(2\pi)^{3}}(K_{1}\cdot K_{2})^{2}\frac{1}{K_{1}^{2}}\frac{1}{K_{2}^{2}}$
$=$ $-\frac{c}{M^{4}}T^{2}\sum_{\omega_{n}\omega_{n}^{\prime}}\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\frac{d^{3}\mathbf{k^{\prime}}}{(2\pi)^{3}}\cdot\mathbf{1}$
$-2\frac{c}{M^{4}}T^{2}\!\!\sum_{\omega_{n}\omega_{n}^{\prime}}\!\!\int\!\!\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\!\frac{d^{3}\mathbf{k^{\prime}}}{(2\pi)^{3}}\!(K\!\cdot\!K^{\prime})^{2}\frac{1}{\omega_{n}^{2}+k^{2}}\frac{1}{\omega_{n}^{\prime 2}+k^{\prime 2}}.$ (35)
where we have relabeled the four external momenta into two independent momenta: $K=(\omega_{n},\mathbf{k})$ and $K^{\prime}=(\omega_{n}^{\prime},\mathbf{k^{\prime}})$ as can be seen in FIG. (1(b)): the diagrams have two independent momenta. And we can drop the first term in Eq. (III), since it vanishes under dimensional regularization. This follows from
$\int d^{d}\mathbf{k}k^{a}$ $=\Omega_{d}\int^{\Lambda}_{0}dkk^{d+a-1}+\Omega_{d}\int^{\infty}_{\Lambda}dkk^{d+a-1}$
$=\Omega_{d}\left(\frac{\Lambda^{3+a-\epsilon_{IR}}}{3+a-\epsilon_{IR}}\right)+\Omega_{d}\left(-\frac{\Lambda^{3+a-\epsilon_{UV}}}{3+a-\epsilon_{UV}}\right),$ (36)
where we written $d=3-\epsilon_{IR}$ for the first integral, assuming $\epsilon_{IR}<0$, and $d=3-\epsilon_{UV}$ for the second integral, assuming $\epsilon_{UV}>0$, because $\epsilon_{IR}$ and $\epsilon_{UV}$ must vanish from physical quantities, we can set $\epsilon_{IR}=\epsilon_{UV}$, then the result of Eq. (III) is zero. Alternatively, we can also simply argue that by dimensional analysis, the integral $\int d^{3}\mathbf{k}\sum_{\omega_{n}}\mathbf{1}$ can only depend on mass $m^{3}$, and in there we consider the $m\to 0$, so the result of this integration will be zero. Thus, we only need to keep the second term:
$f_{(1)}\!=$ $-\!2\frac{c}{M^{4}}T^{2}\!\!\sum_{\omega_{n}\omega_{n}^{\prime}}\!\int\!\!\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\!\frac{d^{3}\mathbf{k^{\prime}}}{(2\pi)^{3}}(K\!\cdot\!K^{\prime})^{2}\!\frac{1}{\omega_{n}^{2}+k^{2}}\!\frac{1}{\omega_{n}^{\prime 2}+k^{\prime 2}}$
$=$ $-\!2\frac{c}{M^{4}}T^{2}\!\!\int\!\frac{d^{3}\mathbf{k}d^{3}\mathbf{k^{\prime}}}{(2\pi)^{6}}\!\!\sum_{\omega_{n},\omega_{n}^{\prime}}\!\frac{\omega_{n}^{2}\omega_{n}^{\prime 2}+2\omega_{n}\omega^{\prime}_{n}kk^{\prime}+k^{2}k^{\prime 2}}{(\omega_{n}^{2}+k^{2})(\omega_{n}^{\prime 2}+k^{\prime 2})}$
$=$ $-\!2\frac{c}{M^{4}}T^{2}\!\!\int\!\frac{d^{3}\mathbf{k}d^{3}\mathbf{k^{\prime}}}{(2\pi)^{6}}\!\!\sum_{\omega_{n},\omega_{n}^{\prime}}\!\frac{\omega_{n}^{2}\omega_{n}^{\prime 2}+k^{2}k^{\prime 2}}{(\omega_{n}^{2}+k^{2})(\omega_{n}^{\prime 2}+k^{\prime 2})}.$ (37)
We drop the middle cross term of the second line due to it is odd form under the integral and will be zero. We now treat the two contributions in the numerator separately, and first calculate the sum and then the integral. For the first part of the numerator we have
$\sum_{w_{n},w_{n}^{\prime}}\frac{\omega_{n}^{2}}{(\omega_{n}^{2}+k^{2})}\frac{\omega_{n}^{\prime 2}}{(\omega_{n}^{\prime 2}+k^{\prime 2})},$
$=\sum_{w_{n},w_{n}^{\prime}}\left[1-\frac{k^{2}}{(\omega_{n}^{2}+k^{2})}\right]\left[1-\frac{k^{\prime 2}}{(\omega_{n}^{\prime 2}+k^{\prime 2})}\right]$
$=\sum_{w_{n},w_{n}^{\prime}}\mathbf{1}-\sum_{w_{n},w_{n}^{\prime}}\mathbf{1}\cdot\frac{k^{2}}{\omega_{n}^{2}+k^{2}}$
$-\sum_{w_{n},w_{n}^{\prime}}\mathbf{1}\cdot\frac{k^{\prime 2}}{\omega_{n}^{\prime 2}+k^{\prime 2}}+\sum_{w_{n},w_{n}^{\prime}}\frac{k^{2}}{\omega_{n}^{2}+k^{2}}\frac{k^{\prime 2}}{\omega_{n}^{\prime 2}+k^{\prime 2}},$ (38)
in which the first three terms vanish after dimensional regularization, for the same reason as earlier. The only nonzero contribution comes from the last part of Eq. (III), and the result is
$\frac{k^{2}}{2kT}\left[1+2n_{B}(|k|)\right]\frac{k^{\prime 2}}{2k^{\prime}T}\left[1+2n_{B}(|k^{\prime}|)\right],$ (39)
by using the standard Matsubara sum, where the $n_{B}(E_{k})\equiv\frac{1}{exp(\beta E_{k})-1}$ is the Bose distribution. For the second part of the numerator we can get:
$\sum_{w_{n},w_{n}^{\prime}}\frac{k^{2}}{(\omega_{n}^{2}+k^{2})}\frac{k^{\prime 2}}{(\omega_{n}^{\prime 2}+k^{\prime 2})}$
$\Rightarrow\frac{k^{2}}{2kT}\left[1+2n_{B}(|k|)\right]\frac{k^{\prime 2}}{2k^{\prime}T}\left[1+2n_{B}(|k^{\prime}|)\right].$ (40)
Then we get the result of Eq. (III)
$-2\frac{c}{M^{4}}T^{2}\int\frac{d^{3}\mathbf{k}\cdot d^{3}\mathbf{{k^{\prime}}}}{(2\pi)^{6}}\left(\frac{|k||k|^{\prime}}{4T^{2}}+\frac{|k||k|^{\prime}}{4T^{2}}\right)$
$\times\left[1+2n_{B}(|k|)\right]\left[1+2n_{B}(|k^{\prime}|)\right].$ (41)
So we have
$-\frac{c}{M^{4}}T^{2}\int\frac{d^{3}\mathbf{k}\cdot d^{3}\mathbf{{k^{\prime}}}}{(2\pi)^{6}}\frac{|k||k^{\prime}|}{T^{2}}$
$\times\left[1+2\frac{1}{exp(\beta|k|-1)}\right]\left[1+2\frac{1}{exp(\beta|k^{\prime}|-1)}\right].$ (42)
Then by the same argument of Eq. (III), we only need to consider the integration
$-\frac{c}{M^{4}}\int\frac{d^{3}\mathbf{k}\cdot d^{3}\mathbf{{k^{\prime}}}}{(2\pi)^{6}}\frac{4|k||k^{\prime}|}{\left[exp(\beta|k|)-1\right]\left[exp(\beta|k^{\prime}|-1)\right]}$
$=-\frac{c}{M^{4}\pi^{4}}\int dk\cdot dk^{\prime}\frac{k^{3}\cdot k^{\prime 3}}{\left[exp(\beta|k|)-1\right]\left[exp(\beta|k^{\prime}|-1)\right]}$
$=-\frac{c\pi^{4}}{225M^{4}\beta^{8}}.$ (43)
Finally, the free energy density, including both the leading-order and first-order contributions, is given by
$f=f_{(0)}+f_{(1)}=-\frac{\pi^{2}T^{4}}{90}-\frac{c\pi^{4}T^{8}}{225M^{4}}.$ (44)
At this point, having incorporated the leading and first-order terms, one might naturally ask whether higher-order corrections, such as the second-order term $f_{(2)}$, should also be considered. To assess this, we estimate the magnitude of subsequent terms in the perturbative expansion. For example,
$f_{(2)}$ $\sim\frac{c^{2}}{M^{8}}T^{12}\sim c^{2}\left(\frac{T}{M}\right)^{8}T^{4},$ (45)
$f_{(3)}$ $\sim\frac{c^{3}}{M^{12}}T^{16}\sim c^{3}\left(\frac{T}{M}\right)^{12}T^{4},$ (46)
and so on. Each term is increasingly suppressed by higher powers of the expansion parameter $(T/M)^{4}$, whether arising from higher-loop corrections or higher-dimensional operators.
Even when infrared divergences emerge at higher loop order—requiring resummation, for instance via thermal mass corrections—the dominant resummed contribution scales as
$f_{\text{resum}}\sim\left(\frac{cT^{4}}{M^{4}}\right)^{3/2}T^{4}\sim c^{3/2}\left(\frac{T}{M}\right)^{6}T^{4},$ (47)
which remains parametrically suppressed for $T\ll M$. Therefore, all higher-order corrections, including resummed IR-sensitive contributions, are negligible within the regime of validity of our analysis.
The entropy density is then given by
$s=-\frac{\partial f}{\partial T}=\frac{2\pi^{2}T^{3}}{45}+\frac{8c\pi^{4}T^{7}}{225M^{4}}.$ (48)
It must be larger than the thermal entropy arising solely from the free part of the theory that does not interact with the heavy particles, from which we immediately deduce that
$c>0.$ (49)
IV the connection to causality and unitarity
Previous work on positivity bounds has primarily focused on analyzing $2\rightarrow 2$ scattering amplitudes to extract constraints from fundamental principles such as causality and unitarity. Causality requires that the scattering amplitude be analytic in the upper half of the complex energy plane, while unitarity implies the optical theorem, which states that the imaginary part of the forward scattering amplitude must be non-negative. By combining these principles within the framework of dispersion relations, one can derive positivity conditions on low-energy parameters of the EFT, such as Wilson coefficients. These coefficients appear in the derivative expansion of the EFT Lagrangian, and positivity bounds require certain combinations of them to be strictly positive.
In our context, unitarity is reflected in the preservation of the trace of the density matrix under unitary time evolution. That is, the total probability of all physical outcomes must remain normalized: $\frac{d}{dt}\mathrm{Tr}(\rho)=0.$ This property underlies the validity of entropy-related inequalities such as the Araki-Lieb inequality.
Causality, while more subtle, is encoded in the $i\epsilon$ prescription of QFT, which ensures microcausality—the requirement that fields commute (for bosons) or anticommute (for fermions) at spacelike separation. Specifically, for a scalar field:
$\phi(x),\phi(y)]=0\quad\text{for}\quad(x-y)^{2}>0.$
This condition guarantees that events separated by spacelike intervals cannot influence each other, consistent with special relativity. This microcausality condition arises from the analytic structure of the Feynman propagator:
$G_{F}(x-y)=\int\frac{d^{4}p}{(2\pi)^{4}}\frac{ie^{-ip\cdot(x-y)}}{p^{2}-m^{2}+i\epsilon},$
where the $i\epsilon$ shifts the poles away from the real axis. The field commutator can be written as the difference between the Feynman and Dyson (time-reversed) propagators:
$\phi(x),\phi(y)]=G_{F}(x-y)-G_{F}(y-x),$
and this vanishes for spacelike separations due to the symmetry of contour deformation in the complex $p^{0}$ plane, enforced by the $i\epsilon$ prescription.
The $i\epsilon$ prescription can also be understood as a regularization of the path integral. One may equivalently insert a term $i\epsilon\phi^{2}(x)$ into the action to ensure convergence for large field amplitudes: $S\rightarrow S-i\epsilon\int d^{4}x\,\phi^{2}(x)$. When performing a Wick rotation to Euclidean space, the prescription dictates that the time variable be rotated anticlockwise in the complex plane: $\tau\rightarrow it$, which ensures that poles in the complex energy plane are avoided appropriately. As a result, both causality and unitarity are preserved in thermal field theory, and they underlie the entropy-based positivity bounds derived in this work.
V Conclusion and discussion
In this work, we computed the thermal entropy of a scalar field theory with a shift symmetry and demonstrated that the requirement of entropy positivity imposes a nontrivial constraint on the sign of higher-dimensional operators in the low-energy effective field theory. In particular, we showed that the coefficient of the leading dimension-8 operator must be strictly positive.
Our derivation was based entirely on thermodynamic consistency, rather than relying on scattering amplitudes or dispersion relations. This offers a complementary perspective on EFT constraints, highlighting the role of entropy and the thermal partition function in encoding information about underlying physical principles such as unitarity and causality.
The thermal entropy approach developed here can be generalized in several directions. For example, it would be natural to extend our analysis to include fermionic or gauge degrees of freedom, or to explore the implications of gravitational interactions. In particular, effective theories involving gravity often pose challenges to traditional positivity bounds due to the presence of massless poles and the absence of an S-matrix in asymptotically de Sitter or AdS spacetimes. Since our approach does not rely on the flat-space S-matrix, it may offer a path forward in these contexts.
Acknowledgements
We are grateful to Anna Tokareva for carefully reading the manuscript and for insightful comments. Y.X. also thanks Longqi Shao and Shuang-Yong Zhou for inspiring discussions.